A piece of metal was found to have 90% purity of iron Fe . An excess of Hcl was added to it . What would be the mass of the piece metal if the volume of H2 produced from this reaction was 14.4 L . Given : Vm=24l.mol


1
Expert's answer
2021-03-08T05:53:59-0500

Fe + 2HCl = FeCl2 + H2

n (H2) = V/22.4 = 14.4/24 = 0.64 mol

n (H2) = n (Fe) = 0.64 mol

n (Fe) = m/M

m (Fe) = n x M

M (Fe) = 55.85 g/mol

m (Fe) = 0.64 x 55.85 = 35.74 g

With 90% purity, a piece of metal must have the mass of:

m (metal) = (35.74 x 100) / 90 = 39.71 g